OpenStudy (anonymous):

Give the complete factorization for the polynomial. 10x - 3 - 3x2 can somebody please explain how to do this???

OpenStudy (anonymous):

So the polynomial can be rearranged as: \[-3x^2+10x-3=0\] note that the polynomial is in the form: \[ax^2+bx+c=0\] To start factoring, find two numbers that multiply to (C*A) So find two numbers that multiply to 9 and add to +10. What are they?

OpenStudy (anonymous):

Could I include negatives?

OpenStudy (anonymous):

You could, but you don't need them here, since both numbers are positive.

OpenStudy (anonymous):

can't 9 only be divided by 1,3 and itself?

OpenStudy (anonymous):

Good, going on that in reverse, what are the only two sets of numbers that multiply to 9? (Two sets of numbers)

OpenStudy (anonymous):

9, 1 and 3, 3. So that would be 1,9 because they equal 10

OpenStudy (anonymous):

Yup, you got it! Now, you know how we have 10x? We're gonna split that up into 9x+1x (it's the same thing) \[-3x^2+9x+1x-3=0\] Now we factor out the negative in the front to make stuff easier for us. \[-(3x^2-9x-1x+3)=0\] Now we want to factor this by groups, the first group is: \[3x^2-9x\] and second group is: \[-1x+3\] So now factor these two grouping individually please.

OpenStudy (anonymous):

@craftshark98, so any ideas how to do it?

OpenStudy (anonymous):

Yeah I'm trying to figure it out right now. Sorry it's taking so long.

OpenStudy (anonymous):

Here's a hint, take out 3x for the first one, and take out -1 for the second one ;D

OpenStudy (anonymous):

what do you mean? :/

OpenStudy (anonymous):

yeah I think I'm pretty stuck lol

OpenStudy (anonymous):

So basically what we do for the first group is take out 3x, since that's the common factor: \[3x^2-9x \rightarrow 3x(x-3)\] And the second group we take out -1, which is the common factor to get: \[-1x+3\rightarrow -1(x-3)\]

OpenStudy (anonymous):

So understand what I did there?

OpenStudy (anonymous):

yeah....but wha I don't understand is how in the second one +3 becomes -3

OpenStudy (anonymous):

So what happens when we divide a + by a negative? It becomes negative right? So that's what I did here.

OpenStudy (anonymous):

Essentially the main idea is to get the same thing in the parenthesis for both groups.

OpenStudy (anonymous):

what did you divide it by?

OpenStudy (anonymous):

-1, so +3/-1 = -3 right?

OpenStudy (anonymous):

yes it is :D but why do we want (x - 3) in the parenthesis of both?

OpenStudy (anonymous):

Since that is one of our factors :D Better move on: So first we have this: \[−(3x^2−9x−1x+3)=0\]Now sub for each group: \[−(3x(x-3)−1(x-3))=0\] Now you see that x-3 is in both sets? That's 1 group, and we factor that out now! So we end up being left with this when factoring: \[-((3x-1)(x-3))=0\] Once we have that, we can get rid of outer parenthesis and we are left with factored form: \[-(3x-1)(x-3)=0\]

OpenStudy (anonymous):

Get it?

OpenStudy (anonymous):

yeah. thank you! :D though I might need help in applying this to other problems.

OpenStudy (anonymous):

It's okay, hope this helped.